# Eka-Eval [Uncaptioned image] : An Evaluation Framework for Low-Resource Multilingual Large Language Models

### v4

@@ 6 Conclusion and Future Work
We present Eka-Eval, an open-source unified evaluation framework designed to overcome persistent challenges in accessibility, reproducibility, and multilingual coverage for language model assessment. Our study with eleven participants demonstrates substantial usability improvements, achieving lowest setup time and the highest satisfaction ratings. Eka-Eval enables an interactive UI, rigorous, reproducible evaluation without manual integration overhead. The current release is an ongoing work, with 30 stars and 3 forks on GitHub. Future development will extend the framework to more than 100 tasks, specialized domains, and advanced diagnostic capabilities, establishing Eka-Eval 2.0 as a comprehensive and equitable standard for multilingual language model evaluation.
@@ 7 Limitations
While Eka-Eval supports a wide range of benchmarks and model backends, it does not yet support multi-modal LLMs. Additionally, reproducibility may be affected by changes in external datasets or model versions if not explicitly versioned or cached.
@@ A.2 Prompt Template System
Figure 4: High-level architectural overview of the Eka-Eval framework.
Appendix B Zero-code Web Interface
To democratize access for LLM evaluation, Eka-Eval includes a comprehensive "Zero-code" web interface. This GUI mirrors the capabilities of the CLI serving as a frontend layer as seen in (Figure 4) while adding advanced features for prompt configuration, resource management, and AI-driven analysis making it accessible for both non-technical and technical users.
Figure 5: High-level benchmark selection dashboard categorized by domain (e.g., Math, Code Generation, Indic Benchmarks).
@@ Appendix B Zero-code Web Interface
For advanced control over the evaluation process, Configuration Panel (Figure 6) offers adjustable settings such as temperature, top-p and batch size.It also includes a built-in GPU resource manager, allowing users to allocate specific or multi CUDA for evaluation.
A key advantage of the web interface is the Prompt Customization Engine (Figure 7). It lets users visually edit system prompts, few-shot templates, and variable fields, removing the need to manually modify JSON configuration files.
@@
(Figure 8) shows real time progress during inference including a live console and, once finished, provides a summary of the model’s performance with options to download logs or explore detailed analysis.
Figure 9: AI diagonosis dashboard providing automated textual insights into model strengths and weaknesses based on performance.
@@
uses LLama 3.3-70b to analyze the evaluation logs, generating summaries of the model’s observed strengths (e.g., "Strong logical coherence") and weaknesses (e.g., "Hallucination in low-resource languages") as example shown in (Figure 9).
Figure 10: AI diagnosis of the wrong result of benchmark selected
@@
Finally, the Model Leaderboard (Figure 11) provides a central place to compare models. It gathers results from all local runs and visualizes performance differences with radar charts and bar graphs, helping users choose the best model.
@@ C.1 CLI Demonstration
Figure 14: Interactive visualisation setup in Eka-Eval.
@@
Figure 16: Bar chart visualisation
@@
Appendix D Computation Requirement and Budget

### v5

@@ 6 Conclusion and Future Work
We present Eka-Eval, an open-source unified evaluation framework designed to overcome persistent challenges in accessibility, reproducibility, and low-resource multilingual coverage for language model assessment. Our study with eleven participants demonstrates substantial usability improvements, achieving lowest setup time and the highest satisfaction ratings. Eka-Eval enables an interactive UI, rigorous, reproducible evaluation without manual coding expertise. The current release is an ongoing work, with 30 stars and 3 forks on GitHub. Future work will expand coverage to multimodal LLMs, benchmarks, and AI diagnostic capabilities.
@@ 7 Limitations
Multimodal Support.
Eka-Eval currently supports text-only LLMs; multimodal model evaluation is left to future work.
Reproducibility.
Evaluation results may be affected by changes in external dataset or model versions; we recommend explicit versioning and caching for fully reproducible runs.
User study scope.
The usability study (n=11) constitutes a formative evaluation conducted under GPU resource constraints, providing initial usability insights and motivating broader future evaluations.
Inference Backend.
Eka-Eval currently relies on HuggingFace Transformers for model inference; vLLM backend integration is underway to enable faster, high-throughput evaluation at scale.
Scalability.
The current release processes requests sequentially; supporting concurrent multi-user deployments may require additional infrastructure and compute beyond the scope of this work.
@@ Appendix A Appendix
   Framework    Winogrande    PIQA    ARC-IN    MMLU-IN
   lm-eval-harness    65.8 $\pm$ 6.0    74.8 $\pm$ 3.5    –    –
   OpenCompass    61.1 $\pm$ 5.0    73.3 $\pm$ 4.3    –    –
   indic-eval    65.8 $\pm$ 1.9    78.7 $\pm$ 1.4    28.1 $\pm$ 2.3    35.0 $\pm$ 2.1
   FreeEval    60.0 $\pm$ 4.5    54.5 $\pm$ 3.2    –    –
   Eka-Eval    66.5 $\pm$ 2.5    77.5 $\pm$ 1.5    31.5 $\pm$ 3.5    33.5 $\pm$ 0.5
@@ A.2 Prompt Template System
Figure 4: The overall architectural pipeline of the Eka-Eval framework.
Appendix B Zero-code User Interface
To democratize access for LLM evaluation, Eka-Eval includes a “Zero-code” user interface that mirrors the full capabilities of the CLI while adding advanced features for prompt configuration, resource management, and AI-assisted analysis, making rigorous evaluation accessible to both non-technical and technical users. Figure 4 illustrates the end-to-end architecture, spanning user input, model inference, evaluation execution, and result visualization across Eka-Eval’s four core components.
Figure 5: Standardized study protocol specifying permitted and prohibited actions for all participants.
Figure 6: Participants assessment form.
Figure 7: ModelHub selection panel to choose any HuggingFace/Local models for benchmark evaluation.
Figure 8: Benchmark selection dashboard categorized by domain (e.g., Math, Code Generation, and other benchmarks.
Figure 12: AI diagnosis dashboard providing automated textual insights into model strengths and weaknesses based on performance.
Figure 13: AI diagnosis showing the wrong result of selected benchmark
Figure 14: Model Leaderboard through aggregated average scores across multiple benchmarks and through live evaluation
Figure 16: Model Key strengths and weaknesses featuring multi-dimensional visualizations along with benchmarks performance analysis for improvement.
@@ Appendix B Zero-code User Interface
For advanced control over the evaluation process, Configuration Panel (as shown in Figure 9) offers adjustable settings such as temperature, top-p and batch size. It also includes a built-in GPU resource manager, allowing users to allocate specific or multi-GPU support for evaluation.
A key advantage of the web user interface is the Prompt Customization Engine (Figure 10). It lets users visually edit system prompts, few-shot templates, and variable fields, GPU allocation, thus removing the need to manually modify JSON configuration files.
@@
Figure 11 shows real time progress during inference including a live console and, once finished, it provides a summary of the model’s performance with options to download logs or explore detailed analysis.
@@
uses LLama 3.3-70b model to analyze the evaluation logs, generating summaries of the model’s observed strengths (e.g., ‘Strong logical coherence’) and weaknesses (e.g., ‘Hallucination in low-resource languages’) as example shown in Figure 12.
@@
Finally, the Model Leaderboard Figure 14, Figure 15 and Figure 16 provides a centralized, dynamic interface for comparative analysis. Unlike static evaluation tables, it aggregates results from all local runs securely. Researchers can filter outputs by model size, language, or specific benchmark, and generate side-by-side multi-dimensional radar charts. To maintain integrity, the leaderboard includes version tracking for datasets and prevents score manipulation by linking directly to the immutable evaluation logs generated by the backend engine.
@@ C.1 CLI Demonstration
Figure 19: Interactive and user-friendly visualisation setup in Eka-Eval.
@@
Figure 21: Bar chart visualisation on Reading Comprehension benchmarks
@@
Appendix D User Study Protocol
Participants received a standardized evaluation protocol and rating form (Figures 5 and 6) to ensure consistency across all framework evaluations. Each participant was permitted to consult official documentation but was prohibited from seeking external technical assistance beyond publicly available framework resources. Explicit guidelines specified permitted and prohibited actions to minimize evaluation bias.
Appendix E Extensibility and Customization
Eka-Eval emphasizes extensibility through a low-code plugin architecture that enables easy integration of new benchmarks and custom metrics with minimal modifications, managed via a hierarchical JSON configuration system that supports complex setups such as parameter sweeps and prompt variations. The framework provides a zero-code UI that enables real-time, interactive adjustment of inference parameters (e.g., temperature, batch size) and evaluation strategies, eliminating the need for users to edit the configuration files directly.
Appendix F Computation Requirement and Budget